OpenStudy (idealist10):

How to integrate (1+(x-1)sec^2 x)/(x-1)?

OpenStudy (amistre64):

split the fraction

OpenStudy (idealist10):

How?

OpenStudy (amistre64):

(a+b)/c = a/c + b/c ....

OpenStudy (amistre64):

or in this case: (a+cb)/c = a/c + b

OpenStudy (idealist10):

I got it.

OpenStudy (amistre64):

\[\int\frac{1}{x-1}+sec^2(x)~dx\]

OpenStudy (idealist10):

Why didn't I thought of that before?

OpenStudy (amistre64):

:)

OpenStudy (idealist10):

:)